1. At 733 mmHg and 114.8oF, calculate the density of HBr gas. Complete solution


1
Expert's answer
2022-02-23T10:27:04-0500

The molar mass of HBr is 80.9 g mol−1

R = gas constant = 0.0821 L atm K−1 mol−1

Converting units:

P = (733 mmHg) × (1 atm / 760 mmHg) = 0.9645 atm

T = (114.8oF + 459.67) × (5/9) = 319.15 K


Solution:

The ideal gas equation can be used.

PV = nRT


Moles (n) = Mass (m) / Molar mass (M)

Therefore, the ideal gas equation can be rewritten as:

PV = (m/M)RT

PVM = mRT


Density (d) = Mass (m) / Volume (V)

Rearrange the equation to make d the subject:

d = m / V = PM / RT

d = PM / RT


Therefore,

Density of HBr gas = (0.9645 atm × 80.9 g mol−1) / (0.0821 L atm K−1 mol−1 × 319.15 K) = 2.978 g/L

Density of HBr gas = 2.978 g/L = 2.98 g/L


Answer: The density of HBr gas is 2.98 g/L